OpenStudy (anonymous):

Find a - 2b + 3c diminished by a - 2b - 3c + d.

OpenStudy (marihelenh):

Diminished is just a fancy way of saying subtract, so you can do the problem the same way as the last one.

OpenStudy (anonymous):

oh okay

OpenStudy (marihelenh):

With this problem, make sure you only subtract like terms.

OpenStudy (anonymous):

-4b + d?

OpenStudy (marihelenh):

a-a=? -2b-(-2b)=? 3c-(-3c)=? 0-d=?

OpenStudy (anonymous):

-4b-6c

OpenStudy (marihelenh):

The b would cancel out because it would become -2b+2b=0. You can add the c together, though. It would be 3c+3c=6

OpenStudy (marihelenh):

0d-d=-1d=d

OpenStudy (marihelenh):

I meant -d for ^^^^

OpenStudy (anonymous):

oh

OpenStudy (marihelenh):

And 6c for the earlier one, sorry. Your final answer should be 6c-d

OpenStudy (anonymous):

oh okay thanks